OpenStudy (anonymous):

the product of two consecutive even integers is 14 less than 7 times their sum. Find the 2 integers

OpenStudy (anonymous):

if one is \(n\) the next is \(n+2\) and the product will be \[n(n+2)\]

OpenStudy (anonymous):

there sum is \(n+n+2=2n+2\) and so 7 times there sum is \(7(2n+2)\) 14 less than that amount is \(7(2n+2)-14\)

OpenStudy (anonymous):

set \[n(n+2)=7(2n+2)-14\] and solve for \(n\)

OpenStudy (anonymous):

so... n^2+2n=15n+14-14 n^2+2n=15n

OpenStudy (anonymous):

n^2+2n-15n=0 n^2-13n=0